OpenStudy (anonymous):

how to find the third degree coefficients of -3 and sqaure root 5

OpenStudy (anonymous):

oops I meant \[\sqrt{2}\]

OpenStudy (phi):

can you add more info... I don't understand the question

OpenStudy (anonymous):

That's all the question says.

OpenStudy (phi):

what are you studying ?

OpenStudy (anonymous):

well it says... Write the simplest third degree polynomial function of -3 and \[\sqrt{2}\]& Algebra 2 Trig

OpenStudy (phi):

that sounds different from "how to find the third degree coefficients of -3 and sqr(2)" but it is still not clear. If it said " Write the simplest third degree polynomial function whose roots are -3 and sqr(2)" we could do that.

OpenStudy (anonymous):

sorry I don't remember exactly,but if you could tell me hoe to do " Write the simplest third degree polynomial function whose roots are -3 and sqr(2)" kind of problems that's be great too.

OpenStudy (phi):

use roots x=-3, x= -\(\sqrt{2}\) and x =\(\sqrt{2}\) x= -3 add +3 to both sides x+3 = 0 x= -\(\sqrt{2}\) add \(\sqrt{2}\) to both sides x+\(\sqrt{2}\) =0 x= \(\sqrt{2}\) add -\(\sqrt{2}\) to both sides x-\(\sqrt{2}\) =0 now multiply all there together (x+3)(x+\(\sqrt{2}\) )(x-\(\sqrt{2}\) )=0 can you multiply that out ?

OpenStudy (anonymous):

I think I can... Thanks!

OpenStudy (anonymous):

@phi I would use the box method, correct?

OpenStudy (phi):

what is the box method ? if you are asking how to multiply out the factors, first do (x + sqr(2))(x- sqr(2)) to get x^2 - 2 now do (x+3)(x^2 - 2) one way is "distribute". in other words, this is the same as A( x^2 - 2)= Ax^2 + A(-2) where A is (x+3)

OpenStudy (anonymous):

sorry, I didn't know you replied. I didn't get a notification,but the box method is when you multiply all the binomials together basically using a box...